### v1

# HTMA-Net: Towards Multiplication-Avoiding Neural Networks via Hadamard Transform and In-Memory Computing

Abstract
Reducing the cost of multiplications is critical for efficient deep neural network deployment, especially in energy-constrained edge devices. In this work, we introduce HTMA-Net, a novel framework that integrates the Hadamard Transform (HT) with multiplication-avoiding (MA) SRAM-based in-memory computing to reduce arithmetic complexity while maintaining accuracy. Unlike prior methods that only target multiplications in convolutional layers or focus solely on in-memory acceleration, HTMA-Net selectively replaces intermediate convolutions with Hybrid Hadamard-based transform layers whose internal convolutions are implemented via multiplication-avoiding in-memory operations. We evaluate HTMA-Net on ResNet-18 using CIFAR-10, CIFAR-100, and Tiny ImageNet, and provide a detailed comparison against regular, MF-only, and HT-only variants. Results show that HTMA-Net eliminates up to 52% of multiplications compared to baseline ResNet-18, ResNet-20, and ResNet-50 models, while achieving comparable accuracy in evaluation and significantly reducing computational complexity and the number of parameters. Our results demonstrate that combining structured Hadamard transform layers with SRAM-based in-memory computing multiplication-avoiding operators is a promising path towards efficient deep learning architectures.
Index Terms—  Deep neural networks, in-memory computing, SRAM, Hadamard transform, multiplication-avoiding
1 Introduction
The rapid growth of deep neural networks (DNNs) has intensified the demand for energy-efficient and high-throughput hardware acceleration [1], [2]. In particular, multiply accumulate (MAC) operations dominate the workload in convolutional neural networks (CNNs), leading to significant energy consumption [3], [4], [5],[6]. Since multiplication is orders of magnitude more costly than addition in digital hardware [7], reducing multiplications is a key strategy for energy-efficient DNN inference. Traditional digital processors struggle to maintain the memory and computation requirements of modern inference workloads, especially in resource-constrained deployment scenarios. This challenge is amplified by the well-known memory wall, where the cost of data movement far exceeds that of arithmetic operations, creating a fundamental bottleneck in both latency and energy consumption.
Compute-in-memory (CIM) architectures have emerged as a promising pathway to alleviate these bottlenecks by performing operations directly within memory arrays [8, 9]. SRAM-based CIM has demonstrated high parallelism and low data movement, making it particularly attractive for inference acceleration. However, conventional CIM implementations still face challenges in mapping multiplication-intensive operations such as convolutions and fully connected (FC) layers. To address this, MF-Net [10, 11] introduced multiplication-avoiding operators with memory-immersed data conversion, showing that rethinking the fundamental building blocks of neural networks can lead to dramatic gains in efficiency.
At the algorithmic level, structured transforms such as the Hadamard transform (HT) have been investigated as low-complexity alternatives to conventional convolution networks, since they replace multiplications with additions and subtractions and reduce the parameters required in the model [12, 13, 14]. However, prior efforts often either (i) replace some convolutional layers with transform-based operators and failed to replace all, or (ii) deploy IMC primitives in isolation without leveraging transform-based redundancy reduction.
Building on this perspective, we propose to integrate the HT layer (HT) with an in-memory multiplication avoiding operator for CIM-based inference. The WHT is an orthogonal linear transform that requires only additions and subtractions, making it inherently hardware- [15]. Its low complexity nature $\mathcal{O}(N\log N)$ and structured mixing properties position it as a compelling candidate to replace conventional layers or to serve as an additional transformation layer that enhances feature representations with minimal hardware overhead. Unlike generic random projections, the deterministic structure of HT maps efficiently to bit-parallel SRAM arrays, aligning well with the strengths of CIM. Within this Hadamard layer, we replace the channel-mixing projections in the convolution block using the multiplication-avoiding operator of MF-Net, thus reducing multiplications in the transform domain.
The rest of the paper is structured as this: Section.2 reviews prior work on in-memory computing and Hadamard-based methods. Section 3 presents our proposed HTMA-Net framework and its integration into ResNet architectures. Section 4 reports our experimental results.
2 Background
In this section, we review the foundational methods that form the basis of our proposed model. We first provide a brief overview of Walsh–Hadamard transform (WHT), which serves as the core building block of our approach and compute-in-memory (CIM) architectures.
Fig. 1: Accuracy vs. normalized MACs across ResNet18 variants (MF, HT, HTMA) for CIFAR-10. Lower multiply accumulate (MAC) indicates fewer multiplications.
2.1 Hadamard Transform
The Walsh-Hadamard Transform (WHT) is an orthogonal linear transform widely used in signal processing and, more recently, in efficient neural network architectures. Let $H_{m}\in\mathbb{R}^{2^{m}\times 2^{m}}$ denote the Hadamard matrix of order $2^{m}$, defined recursively as
$H_{0}=[1],$ (1)
$H_{m}=\frac{1}{\sqrt{2}}\begin{bmatrix}H_{m-1}&H_{m-1}\\ H_{m-1}&-H_{m-1}\end{bmatrix},\quad m>0,$ (2)
which is equivalent to the Kronecker construction
$H_{m}=H_{1}\otimes H_{m-1},\quad H_{1}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}$. (3)
The Hadamard Convolution Theorem states that the dyadic convolution $y=a\ast_{d}x$ of two vectors $a,x\in\mathbb{R}^{N}$ can be expressed as
$\mathrm{WHT}(y)=\mathrm{WHT}(a)\odot\mathrm{WHT}(x),$ (4)
where $\odot$ denotes element-wise multiplication. This result underpins the use of HT layers as efficient substitutes for conventional convolutional filters by replacing spatial-domain multiplications with transform-domain operations [2].
Moreover, Pan et al. [16] introduce an HT-based convolutional layer in which the convolution between input feature maps and kernels is re-parameterized via the WHT for quantum and/or in-memory computing applications.
2.2 Multiplication-Avoiding (MF) In-Memory Operators
In a standard convolutional layer, the output activation is computed as
$y=\sum_{k=1}^{K^{2}C_{\text{in}}}w_{k}\cdot x_{k},$ (5)
where $w_{k}$ are the weights and $x_{k}$ are the activations. This formulation requires multiply–accumulate (MAC) operations, which dominate both energy and latency in deep neural networks.
Table 1: Evaluation comparison of our Hybrid in-memory computing Hadamard Transform (HTMA) based model variants. R = regular conv, HTMA = replaced conv. The evaluation accuracy is based on CIFAR-10 training for 10 epochs.
Layer Regular Middle-only First-only All-stages
Stem Conv R R R R
Stage1-Block1-Conv1 R R R HTMA
Stage1-Block1-Conv2 R R HTMA HTMA
Stage1-Block2-Conv1 R R HTMA HTMA
Stage1-Block2-Conv2 R R HTMA HTMA
Stage2-Block1-Conv1 R R HTMA HTMA
Stage2-Block1-Conv2 R HTMA HTMA HTMA
Stage2-Block2-Conv1 R HTMA HTMA HTMA
Stage2-Block2-Conv2 R R HTMA HTMA
Stage3-Block1-Conv1 R R HTMA HTMA
Stage3-Block1-Conv2 R HTMA HTMA HTMA
Stage3-Block2-Conv1 R HTMA HTMA HTMA
Stage3-Block2-Conv2 R R HTMA HTMA
Stage4-Block1-Conv1 R R HTMA HTMA
Stage4-Block1-Conv2 R R HTMA HTMA
Stage4-Block2-Conv1 R R HTMA HTMA
Stage4-Block2-Conv2 R R HTMA HTMA
Accuracy (%) 85.84 85.94 85.87 85.84
loss 0.4149 0.4152 0.4094 0.4163
MAC 555.4M 404.4M 328.9M 253.4M
MF-Net introduces multiplication-avoiding operators, where multiplications are replaced with additions and bitwise operations, typically mapped onto SRAM-based in-memory computing arrays [10]. Here, MF-Net replaces the scalar product $w_{i}x_{i}$ by an adder-only bilinear-like operator $\oplus$ defined as
$w\oplus x\;\triangleq\;\operatorname{sign}(x)\,|w|\;+\;\operatorname{sign}(w)\,|x|.$ (6)
This identity eliminates the hardware multiplication: it requires only two signed additions on magnitude values plus cheap sign handling. For a vector pair $(\mathbf{w},\mathbf{x})\in\mathbb{R}^{N}$, the MF replacement of the inner product is
$\mathbf{w}\odot_{\mathrm{MF}}\mathbf{x}\;=\;\\ \sum_{i=1}^{N}\!\Big(\operatorname{sign}(x_{i})\,|w_{i}|+\operatorname{sign}(w_{i})\,|x_{i}|\Big)$ (7)
Hence, each multiply–accumulate (MAC) is replaced by two signed additions plus sign logic. In SRAM compute-in-memory (CIM), the magnitudes $|w|$ and $|x|$ can be accumulated via bit-line charge addition, while signs are handled digitally at the periphery, making (7) well aligned with in-memory execution.
3 Methodology
In this section, we describe the Hadamard–Transform convolution layer (HT layer) and its learnable extensions. We then formalize the multiplication-avoiding inner projection and show how it can be embedded within the HT layer.
3.1 Hadamard–Transform Layer
The core contribution of this work is the HTMA block, which combines The Hadamard transform (HT) with multiplication-avoiding (MF) in-memory operators replacing the convolution block. As discussed in Section 2.1, the Hadamard Transform can be integrated into deep neural networks. Similar to [12, 17], we integrate the Hadamard Transform as a learnable layer in different ResNets [2] architectures. Here, given a feature map $\mathbf{X}\!\in\!\mathbb{R}^{H\times W\times C}$, the Hadamard layer operates separably in space using the 2D WHT.
Unlike a pure transform layer, we utilize learnable spectral parameters and projections:
$\mathbf{U}_{i}$ $=\mathbf{U}\odot\mathbf{V}_{i},\qquad i=1,\dots,P$ , (8)
$\mathbf{Z}_{i}$ $=\mathrm{Conv}^{1\times 1}_{i}(\mathbf{U}_{i})$ , (9)
$\hat{\mathbf{Z}}_{i}$ $=\mathrm{SoftThresh}_{\tau}(\mathbf{Z}_{i})$ , (10)
where Eq.8 is per-pod Hadamard mask, Eq.9 is the learnable channel mixing, Eq.10 is sparsity via soft-thresholding, $\mathbf{V}_{i}\!\in\!\mathbb{R}^{H^{\prime}\times W^{\prime}}$ are trainable spectral gates, $P$ is the number of pods, and
$\mathrm{SoftThresh}_{\tau}(u)\;=\;\mathrm{sign}(u)\cdot\max(|u|-\tau,\,0).$
An inverse 2D WHT maps back to the spatial domain, followed by cropping and residual addition. This parameterization turns the HT layer into a learnable spectral convolution, not merely a fixed transform, consistent with the Hadamard–convolutional formulations in [16].
Within this Hadamard Transform layer, the $1{\times}1$ channel-mixing projections (Eq. 9) are implemented using the multiplication-avoiding operator of MF-Net (Eq. 6), thus reducing multiplications in the transform domain. This replacement strategy ensures compatibility with the ResNet design while enabling a significant reduction in MAC operations.
Table 2: CIFAR-10: Accuracy vs. compute between the baseline of ResNet-18 and MF-only, HT-only, and hybrid HTMA variants. ”Rem. MACs” is the number of multiply accumulate, where the letter ”M” refers to Million. “Mult. elim.” is the fraction of baseline multiplications removed.
Model Params (M) Rem. MACs (M) Mult. elim. (%)
ResNet-18 (baseline) 11.1 555 0
ResNet-18 + MF 11.1 8.06 98.5
ResNet-18 + HT 9.86 287 48.3
ResNet-18 + MF + HT (HTMA) 9.86 254 54.4
ResNet-20 (baseline) 0.27 40.8 0
ResNet-20 + MF 0.27 0.71 98.2
ResNet-20 + HT 0.151 21.9 46.2
ResNet-20 + MF + HT (HTMA) 0.151 19.6 52.0
ResNet-50 (baseline) 0.47 69.1 0
ResNet-32 + MF 0.47 0.71 98.9
ResNet-32 + HT 0.265 37.6 45.5
ResNet-32 + MF + HT (HTMA) 0.265 33.7 51.2
Table 3: Generalization on CIFAR-100 and Tiny-ImageNet. Models were trained for 10 epochs.
Acc (%)
Model CIFAR100 Tiny ImageNet
ResNet-18 (baseline) 58.6 49.4
ResNet-18 + MF 35.9 32.3
ResNet-18 + HT 61.8 53.3
ResNet-18 + HTMA 61.1 49.0
3.2 Network Architecture Integration
As summarized in Table 1, we integrate the proposed HTMA blocks into standard residual backbones. We primarily instantiate HTMA within ResNet-18, but also validate on ResNet-20 and ResNet-50 to demonstrate scalability across network depths. Three integration strategies are explored: (i) Middle-only: Only the intermediate residual stages are replaced with HTMA blocks, while others are kept regular. (ii) First-only: The initial stage is kept regular to stabilize low-level representations, whereas all subsequent stages are HTMA. (iii) All: Every residual block is replaced with HTMA, yielding the maximum elimination of multiplications except the Stem convolution block.
4 Results
In this section, we present the results of our proposed HTMA-Net, comparing it against baseline and variant networks. We analyze both arithmetic complexity and classification performance across datasets and architectures.
4.1 Experimental Setup
We evaluate HTMA-Net on three benchmark datasets: CIFAR-10, CIFAR-100, and a subset of Tiny-ImageNet. All models are trained from scratch using standard SGD with momentum, cosine learning rate decay, and standard data augmentation (random cropping, horizontal flipping). ResNet-18 serves as the primary baseline, with ResNet-20 and ResNet-50 included to study scalability.
While our HTMA layers can be implemented in other models as MobileNet [18] or Transformers-based models [19], ResNets are chosen as the backbone family due to their ubiquity in image recognition benchmarks and their modular residual block design, which makes them a natural testbed for layer-level modifications such as MF, HT, and HTMA.
4.2 Experimental Results
Table 1 reports the effect of replacing convolutional layers with HTMA blocks in ResNet-18 on CIFAR-10. The middle-only configuration achieves the best accuracy, slightly surpassing the baseline. Meanwhile, the all-stages configuration reduces multiplications by 54% while maintaining accuracy within 1% of baseline performance. This confirms that HTMA can significantly reduce complexity while preserving accuracy.
Table 2 extends the analysis to different input resolutions and deeper networks (ResNet-20 and ResNet-50). We compare the baseline, MF-only, HT-only, and hybrid HTMA variants under the all-stages integration. Although MF-only yields the largest multiplication reduction, it suffers severe accuracy degradation as Fig. 1). By contrast, HTMA achieves $\geq 51\%$ multiplication elimination with negligible accuracy loss, consistently outperforming MF-only and matching or exceeding HT-only.
Finally, Table 3 summarizes results on CIFAR-100 and Tiny-ImageNet. HTMA outperforms the baseline ResNet-18 on both datasets, highlighting its ability to generalize beyond CIFAR-10 based on the All-Stages method. These results demonstrate that the proposed combination of Hadamard transforms with multiplication-avoiding in-memory computing offers an attractive accuracy–efficiency trade-off for deployment in constrained environments.
5 Conclusion
In this work, we introduced HTMA-Net, a hybrid framework that combines the Hadamard transform with multiplication-avoiding (MA) in-memory computing for efficient convolutional neural networks. This design eliminates up to 52% of multiplications while maintaining accuracy within 1% of the baseline across CIFAR-10, CIFAR-100, and Tiny-ImageNet. HTMA-Net provides a principled way to jointly exploit orthogonal transforms and SRAM-based in-memory computing. This opens a promising direction toward highly efficient deep learning models.

### v2

@@
 We evaluate HTMA-Net on ResNet-18 using CIFAR-10, CIFAR-100, and Tiny ImageNet, and provide a detailed comparison against regular, MF-only, and HT-only variants.
-Results show that HTMA-Net eliminates up to 52% of multiplications compared to baseline ResNet-18, ResNet-20, and ResNet-50 models, while achieving comparable accuracy in evaluation and significantly reducing computational complexity and the number of parameters.
+Results show that HTMA-Net eliminates up to 52% of multiplications compared to baseline ResNet-18, ResNet-20, and ResNet-32 models, while achieving comparable accuracy in evaluation and significantly reducing computational complexity and the number of parameters.
 Our results demonstrate that combining structured Hadamard transform layers with SRAM-based in-memory computing multiplication-avoiding operators is a promising path towards efficient deep learning architectures.
@@
 1 Introduction
-The rapid growth of deep neural networks (DNNs) has intensified the demand for energy-efficient and high-throughput hardware acceleration [1], [2].
-In particular, multiply accumulate (MAC) operations dominate the workload in convolutional neural networks (CNNs), leading to significant energy consumption [3], [4], [5],[6].
+The rapid growth of deep neural networks (DNNs) has intensified the demand for energy-efficient and high-throughput hardware acceleration [1, 2].
+In particular, multiply accumulate (MAC) operations dominate the workload in convolutional neural networks (CNNs), leading to significant energy consumption [3, 4, 5, 6].
 Since multiplication is orders of magnitude more costly than addition in digital hardware [7], reducing multiplications is a key strategy for energy-efficient DNN inference.
@@
 The Walsh-Hadamard Transform (WHT) is an orthogonal linear transform widely used in signal processing and, more recently, in efficient neural network architectures.
+As shown in Fig.2(a), the WHT folow a similar structure as Fast Fourier transform (FFT) and Wavelets [16, 17, 18, 19].
+(a) WHT
+(b) Reduced WHT
+Fig. 2: a) Flow diagram of the N=8 fast Walsh-Hadamard transform, b) upper diagonal of the Walsh-Hadamard flow diagram.
 Let $H_{m}\in\mathbb{R}^{2^{m}\times 2^{m}}$ denote the Hadamard matrix of order $2^{m}$, defined recursively as
@@
 Stage4-Block2-Conv2 R R HTMA HTMA
-Accuracy (%) 85.84 85.94 85.87 85.84
+Accuracy (%) 85.96 85.94 85.87 85.84
 loss 0.4149 0.4152 0.4094 0.4163
@@
 This replacement strategy ensures compatibility with the ResNet design while enabling a significant reduction in MAC operations.
+As a direction for future work, we plan to explore reduced or partial WHT variants, as illustrated in Fig.
+2(b), which may further decrease arithmetic complexity within the transform stage while maintaining performance.
 Table 2: CIFAR-10: Accuracy vs.
@@
 ResNet-20 + MF + HT (HTMA) 0.151 19.6 52.0
-ResNet-50 (baseline) 0.47 69.1 0
+ResNet-32 (baseline) 0.47 69.1 0
 ResNet-32 + MF 0.47 0.71 98.9
@@
 All models are trained from scratch using standard SGD with momentum, cosine learning rate decay, and standard data augmentation (random cropping, horizontal flipping).
-ResNet-18 serves as the primary baseline, with ResNet-20 and ResNet-50 included to study scalability.
+ResNet-18 serves as the primary baseline, with ResNet-20 and ResNet-32 included to study scalability.
 While our HTMA layers can be implemented in other models as MobileNet [22] or Transformers-based models [23], ResNets are chosen as the backbone family due to their ubiquity in image recognition benchmarks and their modular residual block design, which makes them a natural testbed for layer-level modifications such as MF, HT, and HTMA.
@@
 This confirms that HTMA can significantly reduce complexity while preserving accuracy.
-Table 2 extends the analysis to different input resolutions and deeper networks (ResNet-20 and ResNet-50).
-We compare the baseline, MF-only, HT-only, and hybrid HTMA variants under the all-stages integration.
+Table 2 extends the analysis to different input resolutions and deeper networks (ResNet-20 and ResNet-32).
+We compare the baseline, MF-only, HT-only, and hybrid HTMA variants under the Middle-only stage integration.
 Although MF-only yields the largest multiplication reduction, it suffers severe accuracy degradation as Fig.